# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v1

@@ 3.1 Preliminaries
Figure 2: Overview of MoRA. For each new task, we freeze the ranks learned on previous tasks and introduce $r$ new ranks of updates. Our sparse self-activated mixture-of-ranks framework jointly considers all old and new ranks, adaptively inferring a sparse mixture weight for each rank. Panels (a,c) illustrate MoRA conceptually and (b,d) detail its computation for tasks $t$ and $t+1$, respectively.
3.2 Pre-trained Weight Matrices and Low-Rank Updates as Linear Associative Memory
Large PTMs have been shown to lie on a low “intrinsic dimension” manifold [41, 1]. From this perspective, the pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ can itself be viewed as a linear associative memory [37, 5], having stored key–value pair of memories drawn from vast pre-training data. Specifically, any rank-$r$ update $\Delta\mathbf{W}\;=\;\mathbf{B}\,\mathbf{A}$ can be viewed as inserting or updating $r$ key–value entries into the pre-trained weight matrix. Concretely, let
@@ 3.2 Pre-trained Weight Matrices and Low-Rank Updates as Linear Associative Memory
$\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1}^{\top}&\cdots&\mathbf{A}_{r}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\mathrm{in}}},\quad\mathbf{B}=\begin{bmatrix}\mathbf{B}_{1}&\cdots&\mathbf{B}_{r}\end{bmatrix}\in\mathbb{R}^{d_{\mathrm{out}}\times r},$ (2)
where each $\mathbf{A}_{i}\in\mathbb{R}^{1\times d_{\mathrm{in}}}$ serves as a key and each $\mathbf{B}_{i}\in\mathbb{R}^{d_{\mathrm{out}}\times 1}$ as its corresponding value. Then
$\Delta\mathbf{W}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{i}\,\mathbf{A}_{i}$ (3)
writes the value $\mathbf{B}_{i}$ into the memory slot addressed by key $\mathbf{A}_{i}$. In this associative-memory view, each rank-$i$ component explicitly encodes “what to store” (the value $\mathbf{B}_{i}$) and “how to index it” (the key $\mathbf{A}_{i}$), thereby enriching the base model with newly acquired knowledge. More details are left in Appendix A.
@@ 3.3 Mixture of Ranks Adaptation
To overcome the coarse granularity of MoE-LoRA, we introduce mixture-of-ranks, applying MoE at the rank-level subspace within all LoRA adapters. By treating each rank as an independent expert, our fine-grained mixture framework enables: 1. Reduced redundancy and shared knowledge reuse: encoding common, cross-task features (e.g., generic semantics) can be reused, preventing redundant relearning. 2. Minimized interference and enhanced specialization: focusing exclusively on task-specific patterns, ensuring that new knowledge is learned without disrupting existing representations or interfered with future updates.
For any given input token $x$ from task $t$, we weight each rank’s contribution according to its relevance. We introduce a scalar mixture weight $w_{i}\in\mathbb{R}$ for each rank, yielding
$\Delta\mathbf{W}^{t}=\sum\nolimits_{i=1}^{r_{t}}w_{i}\mathbf{B}_{i}\mathbf{A}_{i},$ (4)
where each rank-1 update is analogous to an independent expert, $r_{t}$ is the total number of ranks learned up to task $t$, and $w_{i}$ modulates whether and how strongly the corresponding rank contributes to the final updated weight matrix. Unlike prior work that treats the entire rank-$r$ update as one expert [85, 72], this formulation enables fine-grained utilization of each rank’s update.
@@ 3.4 Self-Activated Adaptive Sparse Mixture of Ranks
Another key challenge in MoE and continual learning is the routing ambiguity and forgetting. To this end, we explore whether each rank-1 expert can autonomously infer its relevance to the current input, eliminating the need for a separate router and enabling self-activated expert selection.
@@
Self-activated sparse mixture of ranks. Rather than relying on an external router, our framework derives mixture weights for each rank directly from the activations of each rank. Recall from Eq. (3) that each rank-$i$ update is parameterized by a key–value pair $(\mathbf{A}_{i},\mathbf{B}_{i})$. Given an input token $x\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, by treating $\mathbf{A}$ as the set of keys, we compute a raw score for all $r_{t}$ ranks accumulated up to task $t$:
$s_{i}\;=\;\frac{\mathbf{A}_{i}x}{\sqrt{\sum_{j=1}^{r_{t}}\|\mathbf{A}_{j}x\|_{2}^{2}}}\quad\in\mathbb{R},$ (5)
@@
where the $\ell_{2}$–normalization over all $r_{t}$ ranks ensures numerical stability and compensates for variations in activation scales between layers and pre-trained weights. As shown in Table 5, using these raw scores alone matches or even exceeds the performance of a learned router for each rank, showing the feasibility that each rank can infer its own relevance using its activations.
To encourage rank specialization and focus the update on the most relevant ranks, we temperature-scale and normalize the scores:
$w_{i}\;=\;\mathrm{softmax}\!\bigl{(}\tfrac{s}{\tau_{\text{MoRA}}}\bigr{)}_{i},$ (6)
where the softmax is taken over the vector $s=[s_{1},\dots,s_{r_{t}}]$ and $\tau_{\text{MoRA}}$ is a temperature hyperparameter that controls the sharpness of the distribution. To ensure only a sparse, specialized subset of ranks contributes, we introduce an activation budget that caps the number of ranks that can be activated.
@@
Adaptive sparse mixture of ranks. To ensure a stable and sparse activation pattern, we impose a fixed sparse rank-activation budget via top-k masking on the $\ell_{2}$–normalized scores $s_{i}:=\mathrm{TopK}(s,\,k),$ where
$[\mathrm{TopK}(s,k)]_{i}=\begin{cases}s_{i},&\text{if }s_{i}\text{ is among the top }k\text{ entries of }s,\\ -\infty,&\text{otherwise,}\end{cases}$ (7)
ensuring a sparse subset of at most $k$ out of the total $r_{t}$ ranks to be activated for task $t$ during CL.
We further refine the mixture weights by applying a fixed threshold $\delta$ to the $\ell_{2}$–normalized scores $s_{i}$. Concretely, after computing the softmax weights $w_{i}$, we define a binary mask $m_{i}=\mathbbm{1}\{s_{i}\geq\delta\}$, and update each gate via $w_{i}:=m_{i}\,\odot\,w_{i}$, where $\mathbbm{1}\{\cdot\}$ is an indicator function and $\odot$ denotes element-wise multiplication. This mask zeroes out any rank whose normalized activation falls below the threshold, ensuring that only the most relevant ranks contribute, resulting in a sparse mixture of an adaptive number of ranks based on the input.
@@
Incremental mixture of ranks. Unlike methods that add a fixed number of LoRA experts per task [85], MoRA incrementally adds $r$ new ranks for each task, freezes all prior ranks, and adaptively selects a sparse subset of the most relevant ranks during both training and inference. This design seamlessly incorporates new task-specific knowledge into freshly initialized ranks while preserving earlier ranks unchanged to prevent forgetting. Moreover, it facilitates common knowledge reuse by allowing ranks that encode common patterns to be activated in later tasks.
@@
Training Objectives. In our experiments, we only use the model’s standard training objective, without any extra regularization or load-balancing constraints. In practice, we empirically observe that the conventional load-balancing loss widely employed in the Mixture-of-Experts literature [65, 40, 24, 16] is not necessary, thanks to our fine-grained self-activated gating design. Imposing a strict load-balancing constraint might hinder the learning of the key vectors (i.e., matrix $\mathbf{A}$), since each rank should remain free to capture its own characteristic input distribution. Fig. 3 illustrates how each rank learns distinct characteristics.
@@
    Method Aircraft [50] Caltech101 [25] DTD [15] EuroSAT [30] Flowers [56] Food [8] MNIST [19] OxfordPet [57] Cars [38] SUN397 [80] Average
@@
   Zero-shot [60] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF [43] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT [78] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL [63] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL [92] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
   MoE-Adapter [85] 71.0 34.9 19.2 63.0 86.6 20.0 87.2 63.7 58.6 56.0
   RAIL-Primal [81] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   CoDyRA [48] 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
@@
   LwF [43] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT [78] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL [63] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL [92] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
   MoE-Adapter [85] 43.6 77.9 52.1 34.7 75.9 86.3 45.2 87.4 66.6 60.2 63.0
   RAIL-Primal [81] 42.4 89.8 55.7 68.5 84.0 83.3 65.3 85.8 67.9 64.5 70.7
   CoDyRA [48] 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
@@
   LwF [43] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT [78] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL [63] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL [92] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
   MoE-Adapter [85] 43.2 78.7 57.6 32.8 79.4 86.0 86.7 87.8 78.2 74.2 70.5
   RAIL-Primal [81] 41.7 94.0 66.0 86.4 97.2 82.4 93.1 83.6 75.0 71.3 79.1
   CoDyRA [48] 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2

### v2

@@ 3.2 Model Weights and Updates as Linear Associative Memory Model
Figure 2: Overview of MoRA. For each new task, we freeze the ranks learned on previous tasks and introduce $r$ new ranks of updates. Our sparse self-activated mixture‐of‐ranks framework jointly considers all old and new ranks, adaptively inferring a sparse mixture weight for each rank. Panels (a,c) illustrate MoRA conceptually and (b,d) detail its computation for tasks $t$ and $t+1$, respectively.
Parameter update as modifications to the linear associative memory. Building on the associative-memory view of weights, we can likewise interpret weight updates during training as key–value insertions and modifications into the existing memory. During training, for an arbitrary update $\Delta\mathbf{W}$ on a given $\mathbf{W}_{0}$ as the starting point, the weight update process can be represented as
$\mathbf{W}_{\text{new}}=\mathbf{W}_{0}+\Delta\mathbf{W},$ (3)
LoRA as low-rank key–value memory updates. LoRA (Hu et al., 2021) performs parameter‐efficient fine‐tuning by adding a rank‐$r$ update $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ to the pre‐trained weight matrix $\mathbf{W}_{0}$, where $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$. With the model update based on LoRA, given an input $\mathbf{x}$, the computation of updated weight becomes
Decomposing the low-rank matrices into a collection of its $r$ rank-one components gives
$\Delta\mathbf{W}\mathbf{x}=\mathbf{B}\mathbf{A}\mathbf{x}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{:,i}(\mathbf{A}_{i,:}\mathbf{x}),$ (6)
@@
with $\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1,:}^{\top}&\cdots&\mathbf{A}_{r,:}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\text{in}}}$ where $\mathbf{A}_{i,:}$ is the $i$-th row of $\mathbf{A}$, and $\mathbf{B}=\begin{bmatrix}\mathbf{B}_{:,1}&\cdots&\mathbf{B}_{:,r}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times r}$ where $\mathbf{B}_{:,i}$ is the $i$-th column of $\mathbf{B}$.
Considering the connection between the formulation of LoRA Eq. (1) and associative‐memory view (Anderson, 1972; Meng et al., 2022) of weight update Eq. (4), we can intuitively interpret LoRA as obtaining a small key–value store in a low‐rank subspace, where the rows of $\mathbf{A}$ serve as keys and the columns of $\mathbf{B}$ serve as their corresponding values. After training the LoRA module on a base weight matrix $\mathbf{W}_{0}$, the modified forward pass in Eq. (5) with Eq. (6) can be seen as: first, the original pre‐trained computation $\mathbf{W}_{0}\,\mathbf{x}$ (Eq. (2)) is performed; then, each learned key $\mathbf{A}_{i,:}$ measures its relevance to the input $\mathbf{x}$ via the product $\mathbf{A}_{i,:}\mathbf{x}$, and each learned value $\mathbf{B}_{:,i}$ is scaled by this relevance score and aggregated, in a same way as shown in Eq. (2). In this way, LoRA effectively inserts or updates a small set of key–value memories that augment the pre‐trained model’s capabilities.
@@ 3.3 Mixture of Ranks Adaptation
In the key–value associative‐memory view of LoRA (Eq. (5) and Eq. (6)), a standard LoRA or its variants (Hu et al., 2021; Meng et al., 2024; Zhang et al., 2024a) apply a dense update: every rank’s subspace contributes to $\Delta\mathbf{W}\,\mathbf{x}$ for any input $\mathbf{x}$. Even in MoE-LoRA, where each LoRA module is treated as a separate expert, all key–value pairs within a LoRA module are typically used densely as a whole (either fully activated or not), which creates redundancy across experts and underutilizes the expressive potential of the key vectors $\mathbf{A}$ to indicate input relevance.
Rather than by default activating all key–value memories for every input, many of which may encode patterns irrelevant to the current token, we introduce a gated, sparse mixture-of-ranks update that selectively activates only the most relevant rank-one key–value pairs. Concretely, we associate each rank-one key–value pair with a scalar weight ${\textnormal{w}}_{i}\in\mathbb{R}$. For an input token $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, with $r_{t}$ total ranks, the update becomes
$\Delta\mathbf{W}^{t}=\sum\nolimits_{i=1}^{r_{t}}{\textnormal{w}}_{i}\mathbf{B}_{:,i}\mathbf{A}_{i,:},{}$ (7)
where ${\textnormal{w}}_{i}$ modulates whether and how strongly each rank $i$ contributes to the updated weight matrix.
@@ 3.4 Self-Activated Adaptive Sparse Mixture of Ranks
A key challenge for MoE-style CL is routing ambiguity, leading to catastrophic forgetting. Conventional routers assign examples to coarse-grained experts, but adding new experts often makes routing suboptimal, causing over-activation or underuse, i.e., the three challenges discussed in Sec. 1. To avoid these pitfalls, we treat each rank-one LoRA (or rank) as an expert based on the rank-level mixture model (Eq. (7)), enabling sparse activation of fine-grained experts for specific tasks. Thus, CL with incrementally added rank-one experts becomes a process of accumulating atomic key–value knowledge that can be reused in future tasks. Unlike conventional MoE-LoRA (Yang et al., 2024a; Wang et al., 2024; Yu et al., 2024), we use the rank-one $\mathbf{A}_{i}$’s (keys) as routers to their corresponding $\mathbf{B}_{i}$’s (values) through self-activation, with each $\mathbf{A}_{i}$$\mathbf{B}_{i}$ pair functioning as a router-expert pair.
@@
Mixture of self-activated rank-one experts. Rather than relying on an external router, our framework derives mixture weights for each rank directly from that rank’s activation. Recall from Eq. (6) that each rank-one update is parameterized by a key–value memory pair $(\mathbf{A}_{i,:},\mathbf{B}_{:,i})_{i=1}^{r_{t}}$; the key $\mathbf{A}_{i,:}$ therefore provides a natural, input-dependent relevance signal. Specifically, given an input token $x\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, and treating $\mathbf{A}$ as the set of keys, we compute a raw score for all $r_{t}$ ranks accumulated up to task $t$:
${\textnormal{s}}_{i}\;=\;\frac{\mathbf{A}_{i}x}{\sqrt{\sum_{j=1}^{r_{t}}\|\mathbf{A}_{j}x\|_{2}^{2}}}\quad\in\mathbb{R},$ (8)
@@
where the $\ell_{2}$-normalization across all $r_{t}$ ranks ensures numerical stability and compensates for activation-scale variations between layers and pre-trained weights. Using raw scores alone matches or surpasses a separately learned router (with 1.5 times parameters) without self-activation (Table 4), showing that each rank can infer its own relevance from activation.
@@
Self-activated rank selection. To ensure a stable and gated, sparse activation pattern, we impose a fixed sparse rank‐activation budget via top-k masking on the $\ell_{2}$–normalized scores ${\textnormal{s}}_{i}:=\mathrm{TopK}({\textnormal{s}},\,k),$ where
[…]
ensuring that at most $k$ out of the total $r_{t}$ ranks are eligible to be activated for task $t$. This encourages rank specialization: only a small set of the most relevant ranks are emphasized and trained to capture each kind of input-specific dynamics, and it prevents tiny, noisy activations from affecting gradients and learning dynamics.
Sharpness control on rank activations. To further encourage rank specialization and concentrate the update on the most relevant ranks, we temperature-scale and normalize the scores:
${\textnormal{w}}_{i}\;=\;\mathrm{softmax}\!\bigl(\tfrac{s}{\tau_{\text{MoRA}}}\bigr)_{i},$ (10)
where the softmax is taken over the vector ${\textnormal{s}}=[{\textnormal{s}}_{1},\dots,{\textnormal{s}}_{r_{t}}]$ and $\tau_{\text{MoRA}}$ controls the distribution’s sharpness. A lower $\tau_{\text{MoRA}}$ concentrates mass on a few ranks (specialists), sharpens the distribution and routes larger learning signals to top ranks, which accelerates specialization and reduces noisy updates to low-relevance ranks, while a higher $\tau_{\text{MoRA}}$ spreads mass (shared representations).
Threshold-based test-time rank selection. To further suppress low-signal ranks at inference, we optionally apply a threshold $\delta$ to the $\ell_{2}$–normalized raw scores and zero out gates below it. We do not hard-threshold during training such that gradients can still flow to near-threshold ranks and allow exploration and adaptation; thresholding only at test time reduces noise and runtime cost:
${\textnormal{w}}_{i}:=\mathbbm{1}\{{\textnormal{s}}_{i}\geq\delta\}\odot{\textnormal{w}}_{i}.$ (11)
This yields a sparse, input-dependent mixture of only the most significant rank-one experts.
Our fine-grained mixture framework offers two key advantages: (1) It reduces redundancy and enables shared-knowledge reuse: ranks that capture common, cross-task features (such as generic semantics) can be reactivated across multiple tasks, preventing those patterns from being relearned redundantly or interfered; (2) It minimizes routing ambiguity and prevents forgetting: each rank can specialize on more nuanced patterns, yielding an optimized, rank‐wise mixture that integrates new knowledge without disrupting existing representations or hindering future updates.
@@
Incremental mixture of ranks. Unlike methods that add a fixed number of LoRA experts per task (Yu et al., 2024), MoRA incrementally adds $r$ new ranks for each task, freezes all prior ranks, and adaptively selects a sparse subset of the most relevant ranks during both training and inference. This design seamlessly incorporates new task-specific knowledge into freshly initialized ranks while preserving earlier ranks unchanged to prevent forgetting. Moreover, it facilitates common knowledge reuse by allowing ranks that encode common patterns to be activated in later tasks.
@@
Training Objectives. In our experiments, we only use the model’s standard training objective, without any extra regularization or load‐balancing constraints. In practice, we empirically observe that the conventional load-balancing loss widely employed in the Mixture-of-Experts literature (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024) is not necessary, thanks to our fine-grained self-activated gating design. Imposing a strict load-balancing constraint might hinder the learning of the key vectors (i.e., matrix $\mathbf{A}$), since each rank should remain free to capture its own characteristic input distribution. Fig. 3 illustrates how each rank learns distinct characteristics.
@@
Method
Aircraft Caltech DTD EuroSAT Flowers Food MNIST OxPet Cars SUN397 Average
@@
   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF (Li & Hoiem, 2017) 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT (Wortsman et al., 2022) 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL (Rebuffi et al., 2017) 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL (Zheng et al., 2023) 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
   MoE-Adapter (Yu et al., 2024) 71.0 34.9 19.2 63.0 86.6 20.0 87.2 63.7 58.6 56.0
   RAIL-Primal (Xu et al., 2024) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   CoDyRA (Lu et al., 2024) 74.3 36.8 44.2 69.9 83.5 42.8 88.9 64.6 63.4 63.2
@@
   LwF (Li & Hoiem, 2017) 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT (Wortsman et al., 2022) 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL (Rebuffi et al., 2017) 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL (Zheng et al., 2023) 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
   MoE-Adapter (Yu et al., 2024) 43.6 77.9 52.1 34.7 75.9 86.3 45.2 87.4 66.6 60.2 63.0
   RAIL-Primal (Xu et al., 2024) 42.4 89.8 55.7 68.5 84.0 83.3 65.3 85.8 67.9 64.5 70.7
   CoDyRA (Lu et al., 2024) 41.4 81.0 58.7 77.8 83.4 84.6 64.5 90.4 67.2 64.4 71.3
@@
   LwF (Li & Hoiem, 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT (Wortsman et al., 2022) 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL (Rebuffi et al., 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL (Zheng et al., 2023) 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
   MoE-Adapter (Yu et al., 2024) 43.2 78.7 57.6 32.8 79.4 86.0 86.7 87.8 78.2 74.2 70.5
   RAIL-Primal (Xu et al., 2024) 41.7 94.0 66.0 86.4 97.2 82.4 93.1 83.6 75.0 71.3 79.1
   CoDyRA (Lu et al., 2024) 37.7 81.5 65.1 89.9 91.4 85.5 96.8 93.3 77.3 73.5 79.2